OpenStudy (anonymous):

write 7*10^-4 as a decimal. please explain how to do this type of problem.

jhonyy9 (jhonyy9):

7*(1/10^4) = 7/(10^4) = 7/10000

jhonyy9 (jhonyy9):

0,00007

jhonyy9 (jhonyy9):

if you need divide 7 by 10000 (10000 there are 4 zeros) so than you need write before 7 four zeros ,like how i have wrote there

OpenStudy (anonymous):

okay, thank you that makes sense i think!

jhonyy9 (jhonyy9):

good luck bye